A reactant decomposes with a half life of 131s with initial M=0.164 M. When initial concentraion=0.768, the same reactant decomposes with a half life of 131s.

What is the order of the reaction and rate constant(k)?
Answer #1

This is a theoretical question, we could use equations to prove them.

Theoretical concept: if less or more amount of reactants is used, there is no change in the rate of reaction, that is, the reactant is NOT a factor in the rate

This is typically for the Zeroth ORder reaction. Rate depends only on the rate constant and not in the concentrations

Therefore,

rate = k

A = A0 - kt

t1/2 = A0/(2k)

The order of reaction is "zeroth order" or simply "0"

the K value

t1/2 = A0/(2k)

K = A0/(2*t1/2) = 0.164 / (2*131) = 0.000625954 M/s